OpenStudy (anonymous):

A ball is thrown into the air by a baby alien on a planet in the system of Alpha Centauri with a velocity of 42 ft/s. Its height in feet after t seconds is given by y=42t−29t2.

hartnn (hartnn):

ok, and find what ?

OpenStudy (anonymous):

A. Find the average velocity for the time period beginning when t=3 and lasting

OpenStudy (anonymous):

.005 s

OpenStudy (anonymous):

For the above answers, you may have to enter 6 or 7 significant digits if you are using a calculator. B. Estimate the instanteneous velocity when t=3.

hartnn (hartnn):

do you have an idea how to start ?

OpenStudy (anonymous):

yes a little bit. avg=displacement/ time elapsed?

hartnn (hartnn):

yes, in terms of differentials, we have the same formula as dy/dt = velocity where y= displ. , t= time. so what you get after differentiating y=42t−29t^2.

hartnn (hartnn):

differentiating y=42t−29t^2. with respect to t

OpenStudy (anonymous):

time last? which is 0.05?

hartnn (hartnn):

differentiate y=42t−29t^2. w.r.t 't' can you ?

OpenStudy (anonymous):

mmm?

OpenStudy (anonymous):

ohh!

OpenStudy (anonymous):

3+0.05?

hartnn (hartnn):

\(d/dt (t^n) = n t^ {n-1} \) so,\( d/dt (y) =d/dt (42t-29t^2)=d/dt(42t)-d/dt(29t^2) \\ =42 d/dt (t)- 29 d/dt (t^2)=...?\)

OpenStudy (anonymous):

oh no im confused

hartnn (hartnn):

say example if y= x^2 dy/dx = 2x right ? (using the formula d/dx (x^n)=n x^{n-1}) now here the variable is 't' instead of x.

hartnn (hartnn):

\(d/dt (t^n) = n t^ {n-1} \\d/dt (t^2) = ....?\)

OpenStudy (anonymous):

nt^2-1!

hartnn (hartnn):

note that n=2 here.

OpenStudy (anonymous):

OK,!

hartnn (hartnn):

so, \(d/dt (t^n) = n t^ {n-1} \\d/dt (t^2) = 2t \\ d/dt (t)=dt/dt =1\) got this much ?

hartnn (hartnn):

doubts ?

OpenStudy (anonymous):

yess.. so far.

OpenStudy (anonymous):

taking notes please hold

hartnn (hartnn):

\(d/dt (y) \\ =d/dt (42t-29t^2)\\= d/dt(42t)-d/dt(29t^2) \\ =42 d/dt (t)- 29 d/dt (t^2)\) look at these steps, and ask if you don't get any particular step...

OpenStudy (anonymous):

Have you actually learnt differentiation in class? @Dodo1

hartnn (hartnn):

\(d/dt (y) \\ =d/dt (42t-29t^2)\\= d/dt(42t)-d/dt(29t^2) \\ =42 d/dt (t)- 29 d/dt (t^2) \\ =42-29 \times (2t) \\ =...?\)

OpenStudy (anonymous):

26t?

hartnn (hartnn):

26? how?

OpenStudy (anonymous):

42-29=13. 13*2t. 2t=13 t=13/2

hartnn (hartnn):

you can't do 42-29 according to PEDMAS, first multiplication needs to be done beffore subtraction.

hartnn (hartnn):

\(d/dt (y) \\ =d/dt (42t-29t^2)\\= d/dt(42t)-d/dt(29t^2) \\ =42 d/dt (t)- 29 d/dt (t^2) \\ =42-29 \times (2t) \\ so, dy/dt =42-58t\) make sure you get/understand all steps of above.

OpenStudy (anonymous):

It's not possible to teach some how to run when they can't even walk yet. That's something you need to think about @Dodo1

OpenStudy (anonymous):

Ok im taking notes please hold

OpenStudy (anonymous):

dear whom it may concern, but you can try walk to get to thepoint of how to.

OpenStudy (anonymous):

I mean run

OpenStudy (anonymous):

Dude, it's a proverb. There's a meaning to that sentence. It means that you have to learn the basics before going ahead.

hartnn (hartnn):

@Azteck can you please continue your comments in messages ? because they are just a distraction from the problem,Thanks :) and atleast i am trying...

OpenStudy (anonymous):

Thank you for your advice people.. I guess i need to start from basic but I need to finish this problem

hartnn (hartnn):

so, everything clear till dy/dt = 42 -58t ?

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

I know you are. But it's really difficult to learn maths when your basics are lacking a little. When someone is not understanding fully, you must know when to stop and advise them to learn the stuff that he/she is confused at. No point continuing further, because it's certain that he/she will not understand.

hartnn (hartnn):

now comes the easy part. you want average velocity at t=3 so, you need to put t=3 in that equation.

OpenStudy (anonymous):

OK so 42-58*3=132

OpenStudy (anonymous):

hold on thatws wrong

OpenStudy (anonymous):

-48

hartnn (hartnn):

you missed the - sign again! 42-58*3 = -132

OpenStudy (anonymous):

Oh. i should use a cal not on comp

OpenStudy (anonymous):

so -132. this is....

hartnn (hartnn):

no...still few steps are there..

hartnn (hartnn):

thats dy/dt = -132

OpenStudy (anonymous):

Ok!

hartnn (hartnn):

avg vel.=displacement/ time elapsed so, you need displacement, and time elapsed given = 0.05 right ?

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

mmm when do I use 3.00?

OpenStudy (anonymous):

I mean 0.005

hartnn (hartnn):

we used dt =0.05 dy = -132 *(0.05) then, avg. velocity = -132 *0.05 / 0.05 =-132...

OpenStudy (anonymous):

mmm so its -132?

hartnn (hartnn):

i think, yes.

hartnn (hartnn):

do you have choices/options ?

OpenStudy (anonymous):

sorry my email box is not working

OpenStudy (anonymous):

No, i dont have any choices

hartnn (hartnn):

then try entering -132..

OpenStudy (anonymous):

it says incorrect

OpenStudy (anonymous):

mm:/

hartnn (hartnn):

i think -132 should work try -132 ft/s or +132 ft/s

hartnn (hartnn):

@Azteck any inputs ?

OpenStudy (anonymous):

but for 0.01 , the answer says 132.29

OpenStudy (anonymous):

-132.29

OpenStudy (anonymous):

 That equation is pretty much the same as \[s=u_{y}t+\frac{1}{2}a_{y}t^2\] What we're finding is the average velocity when t=3 lasting 0.05secs That 0.05 is putting me off.

hartnn (hartnn):

average we find in an interval, right... ? here interval is from 3 to 3.005

OpenStudy (anonymous):

You get -135 when t=3 and then I think you sub 0.05 into the equation for t. And then you either subtract or add.

OpenStudy (anonymous):

That's what I was going to write.

OpenStudy (anonymous):

thank you people sincerely, for helping me tho!

OpenStudy (anonymous):

and taking your time!